# A Survey on Wi-Fi Sensing Generalizability: Taxonomy, Techniques, Datasets, and Future Research Prospects

### v2

@@ II-C4 Meta Learning for Few-shot Learning
Meta learning aims to enable models to “learn how to learn”, and represents a distinct paradigm from metric learning approaches described in Section. II-C3. While both are designed for few-shot learning, metric learning focuses on constructing an embedding space where intra-class samples are pulled closer and inter-class samples are pushed apart, thereby allowing direct inference based on similarity measures. In contrast, meta learning emphasizes learning an optimal initialization or training strategy that can rapidly adapt to new tasks with only a few labeled samples. Model-Agnostic Meta-Learning (MAML) [194] is one of the most representative and widely adopted approaches.
@@
In addition, researchers have explored more recent meta-learning paradigms to improve few-shot cross-domain performance. ML-WiGR [200] incorporates Reptile [267], a first-order approximation to MAML, enabling efficient meta-training without second-order gradients. MetaGanFi [154] adopts the BOIL [268] method, which focuses meta-learning on the feature extractor rather than the task-specific head. Furthermore, Zhang et al. [201] utilize Learning to Learn [269], a meta-optimizer-based approach, to support cross-domain activity recognition.
II-C5 Data Augmentation and Data Synthesize
Data scarcity and limited diversity are key challenges that hinder Wi-Fi sensing generalization. To address these issues, a growing number of studies leverage data augmentation and synthetic data generation. Notably, data generation is often used in conjunction with other methods, such as domain adaptation, metric learning, or meta-learning, serving as a plug-in to enhance overall sensing robustness.
@@ II-C5 Data Augmentation and Data Synthesize
Figure 16: Wi-Fi sensing generalization methods in the deployment stage.
When limited target domain data is available, augmentation strategies can be used to synthesize domain-aligned samples: Wi-Cro [156] uses CycleGAN [253] to generate synthetic samples mimicking the target domain distribution, mitigating domain shift. SFTSeg [174] employs a sliding window strategy to segment Wi-Fi signals in target domain and takes segmented signals to augment training data. OneFi [68] simulates gesture rotations to generate direction-variant samples for cross-orientation recognition. OneSense [92] generates new samples by composing gesture primitives based on signal propagation priors—e.g., combining “L” and “I” to simulate “U”—thereby enabling recognition of unseen gestures. Wang et al. [218] propose a feature decoupling approach that employs two separate classifiers to disentangle identity and gesture features. The model extracts identity-specific and gesture-specific representations independently, and then combines the gesture feature with a target user’s identity representation to generate synthetic gesture samples for that specific user for gesture recognition.
@@ II-D2 Federated Learning
Some works adopt frameworks similar to federated learning but use different terminology. For instance, Co-WiSensing [232] proposes a cloud-edge collaborative framework. In this approach, a cloud server trains a cloud Wi-Fi sensing model using data from various edge devices. The higher layers of the trained model are pruned, and the lower layers are distributed to the edge devices. At the edge, the pruned model is extended with a task-specific head, which is fine-tuned using the edge user’s local data to generate a Wi-Fi sensing model tailored for the edge environment. Zhang et al. [119] also employs a central server and adversarial learning to extract shared attributes across different environments. The activity-oriented attributes are then transferred to the local server, enabling cross-environment Wi-Fi human activity recognition. AdaWiFi [233] proposes a collective sensing framework, where a master device collects data using multiple sensors and trains multiple encoders collaboratively, which are then aggregated into a unified model. In the end-user environment, the model is collaboratively tuned with a small amount of labeled data. Zhang et al. [234] propose a local–global modeling method for indoor localization. The approach involves clustering Wi-Fi fingerprints of a large environment into multiple sub-areas, training local localization models for each sub-area, and then aggregating these local models into a global model. The authors demonstrate that this method achieves superior cross-environment generalization.
@@ III Dataset
Acquiring Wi-Fi sensing data is crucial for effective Wi-Fi sensing applications. Several open-source tools are available for extracting Wi-Fi Channel State Information (CSI) and Beamforming Feedback Information (BFI), as summarized in Table I. Among these, Linux CSI Tool [276] was the earliest open-source CSI extraction tool released in 2011, while PicoScenes currently offers the most advanced CSI extraction capabilities. In addition, due to the limited availability of networks and devices supporting open CSI extraction, BFM-Tool [277] facilitates Beamforming Feedback Information (BFI) extraction, enabling Wi-Fi sensing across a broader range of devices.
TABLE I: Wi-Fi Channel State Information (CSI) and Beamforming Feedback Information (BFI) Extraction Tools
Name
max.
MIMO
802.11
support
max.
# subcarriers
max.
bandwidth
Wi-Fi
Intel CSITool [276] 2$\times$2 n 60 40 MHz CSI
@@
Nexmon CSITool [278] 4$\times$4 n/ac 256 80 MHz CSI
Wi-ESP [279] 2$\times$2 b/g/n/ac 64 40 MHz CSI
ZTE CSITool [280] 3$\times$2 n/ac/ax 512 160 MHz CSI
PicoScenes [281, 282] 4$\times$4 a/g/n/ac/ax/be 1024 320 MHz CSI
BFM-Tool [277] 4$\times$4 ac/ax 512 160 MHz BFI
Over the years, many researchers have generously released public datasets, significantly improving the efficiency of algorithm evaluation by removing the need to collect data from scratch. Among the various sensing tasks, activity recognition is currently the most well-represented in public datasets, as summarized in Table II. To provide a more intuitive overview of activity recognition datasets, Fig. 17 presents a bubble chart that visualizes the number of activity classes, release date, and dataset volume.

### v4

@@ II-C4 Meta Learning for Few-shot Learning
Meta learning aims to enable models to “learn how to learn”, and represents a distinct paradigm from metric learning approaches described in Section II-C3. While both are designed for few-shot learning, metric learning focuses on constructing an embedding space where intra-class samples are pulled closer and inter-class samples are pushed apart, thereby allowing direct inference based on similarity measures. In contrast, meta learning emphasizes learning an optimal initialization or training strategy that can rapidly adapt to new tasks with only a few labeled samples. Model-Agnostic Meta-Learning (MAML) [53] is one of the most representative and widely adopted approaches.
@@
In addition, researchers have explored more recent meta-learning paradigms to improve few-shot cross-domain performance. ML-WiGR [59] incorporates Reptile [152], a first-order approximation to MAML, enabling efficient meta-training without second-order gradients. MetaGanFi [284] adopts the BOIL [155] method, which focuses meta-learning on the feature extractor rather than the task-specific head. Furthermore, Zhang et al. [303] utilize Learning to Learn [6], a meta-optimizer-based approach, to support cross-domain activity recognition. From a networking perspective, meta-learning typically incurs limited communication overhead during deployment, as model adaptation is performed locally using a small number of samples. However, the meta-training phase is often computationally intensive and may introduce higher offline training latency, especially when learning across a large number of tasks or domains.
II-C5 Data Augmentation and Data Synthesis
Data scarcity and limited diversity are key challenges that hinder Wi-Fi sensing generalizability. To address these issues, a growing number of studies leverage data augmentation and synthetic data generation. Notably, data generation is often used in conjunction with other methods, such as domain adaptation, metric learning, or meta-learning, serving as a plug-in to enhance overall sensing robustness.
@@ II-C5 Data Augmentation and Data Synthesis
When limited target domain data is available, augmentation strategies can be used to synthesize domain-aligned samples: Wi-Cro [145] uses CycleGAN [326] to generate synthetic samples mimicking the target domain distribution, mitigating domain shift. DiRA [214] develops a robust data augmentation scheme using conditioned diffusion models [40] to expand both the quantity and quality of wireless sensing data, effectively enhancing the accuracy and robustness of sensing models where data is insufficient or unevenly distributed. SFTSeg [239] employs a sliding window strategy to segment Wi-Fi signals in target domain and takes segmented signals to augment training data. OneFi [244] simulates gesture rotations to generate direction-variant samples for cross-orientation recognition. OneSense [306] generates new samples by composing gesture primitives based on signal propagation priors, e.g., combining the hand-drawing “L” gesture and the hand-drawing “I” gesture to simulate the hand-drawing “U” gesture, thereby enabling recognition of unseen gestures. Wang et al. [216] propose a feature decoupling approach that employs two separate classifiers to disentangle identity and gesture features. The model extracts identity-specific and gesture-specific representations independently, and then combines the gesture feature with a target user’s identity representation to generate synthetic gesture samples for that specific user for gesture recognition.
@@ II-C6 Pseudo Labeling
TABLE IV: Comparison of Pros, Cons, and Best-fit Scenarios of Feature Learning Techniques in Wi-Fi Sensing.
Category Pros (Advantages) Cons (Limitations) Generalization Mechanism Best-fit Scenarios
Domain Alignment ([230, 178, 286, 321, 329, 125, 278, 96, 105, 203, 296, 328, 186, 136, 261, 257, 164, 13, 319, 250, 116, 179, 27, 320, 168, 101, 129, 294, 146, 318, 245, 100, 34, 99, 62, 176, 25, 271, 200, 43, 284, 279, 145, 242, 283, 187, 300, 92, 208]) Mitigates domain and hardware shifts; statistical metrics provide theoretical grounding. Risk of suppressing discriminative motion features during aggressive alignment. Enforces distribution-level invariance by aligning feature statistics across domains (e.g., adversarial learning, MMD). Large domain shifts (e.g., cross-room or cross-device) with labeled source and unlabeled target data.
Disentanglement ([50, 29, 210, 74, 234]) Strong physical interpretability; separates motion features from environment or identity noise. Requires complex architectures (e.g., GANs or autoencoders) and careful hyperparameter tuning. Factorizes latent representations into independent components to isolate task-relevant signals. Scenarios with multiple interference factors (e.g., identical gestures by different users or locations).
Metric Learning ([316, 233, 19, 138, 297, 237, 239, 248, 227, 241, 172, 128, 308, 215, 86, 290, 291, 182, 183, 184, 307, 44, 9, 46, 67, 201, 252, 298]) Native few-shot and zero-shot support; scalable to new classes via template updates. Highly dependent on support set quality; risk of class overlap in the embedding space. Structures the embedding space by minimizing intra-class distance and maximizing inter-class separability. Scenarios with extreme scarcity of target labels or dynamic needs for novel class recognition.
Meta Learning ([90, 322, 56, 177, 306, 52, 42, 59, 284, 303]) Rapid adaptation to new tasks; learns an initialization for fast convergence. High computational cost during offline meta-training (e.g., second-order gradients). Optimizes model parameters for fast task-level adaptation through episodic training across domains. Edge deployment requiring immediate high accuracy after minimal user calibration.
Data Synthesis ([125, 215, 285, 91, 226, 290, 287, 240, 27, 26, 249, 250, 241, 314, 206, 113, 76, 132, 39, 200, 84, 137, 199, 145, 214, 239, 244, 306, 216]) Directly mitigates data scarcity; generative models improve coverage of rare cases. Potential sim-to-real gap between synthetic samples and real-world signal data. Expands training distributions by synthesizing diverse or unseen signal patterns. Plug-in augmentation for tasks with insufficient, imbalanced, or hard-to-collect labeled data.
Pseudo Labeling ([317, 25, 273, 178, 272]) Leverages abundant unlabeled target data via self-training. Risk of error accumulation from early incorrect pseudo-labels. Iteratively refines decision boundaries using confident predictions as surrogate labels. Semi-supervised sensing tasks with long-term accumulation of unlabeled data.
Table IV compares the advantages and limitations of representative feature learning strategies in Wi-Fi sensing, explains their underlying generalization mechanisms, and summarizes the scenarios in which each approach is best suited. The analysis of the feature learning stage highlights several strategic insights, as highlighted in the Key Takeaways and Lessons learned below, for developing robust and cross-domain representations, following the technical progression outlined in Section II-C.
Key Takeaways and Lessons Learned: Feature Learning Domain Alignment for Distribution Matching: Minimizing domain divergence via adversarial discriminators or statistical metrics (e.g., MMD, EMD) is fundamental. A key lesson is using cross-modal embeddings (e.g., BERT, GPT-4o) as “semantic anchors” to bridge the gap between raw CSI and high-level human activities. Component Disentanglement for Interpretability: Explicitly separating environment, identity, and motion components using GANs or specialized encoders ensures that the model ignores environmental noise. This disentangled latent space is critical for handling multi-factor variations in real-world settings. Metric Learning for Few-shot Scalability: Structuring the feature space using triplet loss, contrastive learning, or prototypical networks allows for clustering similar actions. The major takeaway is the high scalability of similarity-based matching, which enables zero/few-shot recognition of novel classes without retraining the backbone. Meta-Learning for Rapid Adaptation: Paradigms like MAML and Reptile focus on “learning how to learn” by optimizing for a superior parameter initialization. This ensures that the sensing system can converge to a site-specific optimum with minimal gradient steps (e.g., one-shot or few-shot). Data Synthesis as a Robustness Plug-in: Beyond simple augmentations, generative models (GANs, VAEs, and Diffusion) are essential for simulating target-domain conditions. These modules mitigate the Sim-to-Real gap by expanding training diversity with physically plausible synthetic signals. Pseudo Labeling for Iterative Refinement: Leveraging abundant unlabeled target data through virtual labels reduces annotation costs. The iterative refinement of these labels (e.g., via confidence thresholds or centroid-based reassignment) is vital for long-term performance improvement in unfamiliar domains.
Figure 16: Wi-Fi sensing generalization methods in the deployment stage. (a) Few-shot fine-tuning. In this approach, a Wi-Fi sensing network is first pre-trained on a large-scale dataset to learn general feature extraction capabilities. During deployment, the parameters of the backbone are kept frozen, while only the task head is fine-tuned using data from the target scenario. This strategy enables rapid deployment in new environments with minimal computational overhead. (b) Federated Learning. This method involves aggregating data from diverse scenarios for centralized training on the server side. By exposing the model to a sufficiently wide range of environments during the training phase, it learns to capture robust, cross-domain features. Consequently, the model can be deployed into target scenarios, achieving seamless zero-shot generalization. (c) Continual Learning. To handle the continuous influx of new sensing tasks and environmental data, the model evolves through sequential training stages as new data arrives, allowing it to incrementally acquire knowledge from novel scenarios. This approach ensures the model remains adaptive to dynamic environmental changes while mitigating the risk of catastrophic forgetting.
@@ II-D Deployment Stage
TABLE V: Comparison of Model Deployment Strategies in Wi-Fi Sensing.
Approach Pros (Advantages) Cons (Limitations) Generalization Mechanism Best-fit Scenarios
Transfer Learning ([208, 113, 43, 30, 17, 51, 260, 310, 14, 82, 83]) Fast deployment by leveraging pre-trained weights.
Resource efficient with minimal fine-tuning on edge devices.
Requires labeled target samples.
Performance degrades under large domain gaps.
Adapts shared representations by fine-tuning higher layers to capture environment- or user-specific patterns. Rapid adaptation to a specific new environment, user, or novel sensing task.
Federated Learning ([78, 60, 126, 160, 293, 286, 311, 282]) Preserves data privacy as raw CSI remains local.
Benefits from collaborative learning across clients.
Communication overhead from frequent model synchronization.
Latency in large-scale deployments.
Learns globally robust models by aggregating heterogeneous local updates from diverse environments. Privacy-sensitive applications such as home security or multi-user collaborative sensing.
Continual Learning ([252, 10, 185, 270, 54, 295, 301]) Incrementally acquires new knowledge with local updates.
Communication efficient for long-term deployment.
Risk of catastrophic forgetting.
Potential inference or update latency on COTS devices.
Maintains generalization over time by balancing knowledge retention and adaptation to new data distributions. Long-term deployments where environments, furniture layouts, or user behaviors evolve.
@@ II-D2 Federated Learning
Some works adopt frameworks similar to federated learning but use different terminology. For instance, Co-WiSensing [293] proposes a cloud-edge collaborative framework. In this approach, a cloud server trains a cloud Wi-Fi sensing model using data from various edge devices. The higher layers of the trained model are pruned, and the lower layers are distributed to the edge devices. At the edge, the pruned model is extended with a task-specific head, which is fine-tuned using the edge user’s local data to generate a Wi-Fi sensing model tailored for the edge environment. Zhang et al. [286] also employs a central server and adversarial learning to extract shared attributes across different environments. The activity-oriented attributes are then transferred to the local server, enabling cross-environment Wi-Fi human activity recognition. AdaWiFi [311] proposes a collective sensing framework, where a central device collects data using multiple sensors and trains multiple encoders collaboratively, which are then aggregated into a unified model. In the end-user environment, the model is collaboratively tuned with a small amount of labeled data. Zhang et al. [282] propose a local–global modeling method for indoor localization. The approach involves clustering Wi-Fi fingerprints of a large environment into multiple sub-areas, training local localization models for each sub-area, and then aggregating these local models into a global model. The authors demonstrate that this method achieves superior cross-environment generalization.
While federated learning effectively preserves data privacy by keeping raw CSI data on local devices, it introduces non-negligible communication overhead due to frequent model parameter exchanges between clients and the central server. The resulting communication latency can become a bottleneck, particularly in large-scale deployments with heterogeneous network conditions and limited uplink bandwidth.
@@ II-D3 Continual Learning
Continual learning primarily relies on local model updates and thus incurs minimal communication overhead during operation. However, frequent online adaptation may introduce additional inference or update latency on resource-constrained devices, highlighting the trade-off between real-time responsiveness and long-term adaptability.
Table V compares the advantages and limitations of representative model deployment strategies in Wi-Fi sensing, explains their underlying generalization mechanisms, and summarizes the scenarios in which each approach is best suited. The analysis of the model deployment stage highlights several strategic insights, as highlighted in the Key Takeaways and Lessons learned below, for developing generalizable Wi-Fi sensing systems.
Key Takeaways and Lessons Learned: Model Deployment Hierarchical Efficiency in Transfer Learning: While fine-tuning the entire network yields superior accuracy, freezing the backbone and updating only the task head is more viable for resource-constrained edge devices. This strategy balances global representation with site-specific adaptation. Privacy-Preserving Collaborative Generalization: Federated Learning is a powerful engine for building robust global models without moving raw CSI data. A major lesson is managing communication overhead during parameter exchange and addressing the performance drops caused by shifted data distributions across heterogeneous clients. Evolution through Continual Learning: To handle environmental drifts (e.g., furniture moving), models must incrementally acquire knowledge. The key challenge is mitigating catastrophic forgetting; strategies like replay and model expansion help preserve performance on original tasks. Minimizing User Burden in Online Adaptation: Incorporating “User-in-the-loop” mechanisms allows for continuous refinement. The key takeaway is using anomaly detection and high-confidence sampling to minimize the labeling effort required from users during the adaptation phase. Systems-level Latency and Resource Trade-offs: Beyond accuracy, real-world deployment on COTS chipsets requires balancing model update frequency with inference latency. For high-responsiveness applications, local updates must be optimized to prevent system bottlenecks.
To further explain why approaches in the four stages of the Wi-Fi sensing pipeline can enhance generalizability, we analyze them through the lens of Domain Adaptation Theory [12]. The generalization error on a target domain $\epsilon_{T}(h)$ is bounded by the source error $\epsilon_{S}(h)$, the domain divergence $d_{\mathcal{H}\Delta\mathcal{H}}(P_{S},P_{T})$, and the complexity of the ideal joint hypothesis $\lambda$:
$\epsilon_{T}(h)\leq\epsilon_{S}(h)+\frac{1}{2}d_{\mathcal{H}\Delta\mathcal{H}}(P_{S},P_{T})+\lambda$
Theoretical Analysis into Generalizability Experimental Setup (Distribution Approximation): By scaling up datasets and utilizing distributed hardware, these approaches ensure the empirical training distribution provides a denser coverage of the physical signal space. This not only minimizes $\epsilon_{S}(h)$ but also constrains the inherent bias $\lambda$ between source and target environments. Signal Preprocessing (Nuisance Variable Elimination): Physics-based indicators (e.g., BVP, DFS) act as canonical coordinate transformations that project raw CSI into a domain-invariant physical space. By decoupling activity features from environmental nuisance variables, these methods significantly reduce the domain divergence $d_{\mathcal{H}\Delta\mathcal{H}}$ at the input level. Feature Learning (Information Bottleneck): Techniques like domain alignment and disentanglement operate on the Information Bottleneck principle [192]. They maximize the mutual information $I(Z;Y)$ for task labels while minimizing $I(Z;D)$ for domain-specific patterns, thereby aligning feature manifolds in a latent subspace. Model Deployment (Bayesian Adaptation): Strategies such as few-shot fine-tuning represent a Bayesian posterior update. They leverage a robust prior $P(\theta)$ learned across diverse source domains to rapidly converge to a site-specific optimum $P(\theta|D_{target})$ with minimal target data.
Highlight While this survey focuses on Wi-Fi sensing, many of the taxonomy and generalization challenges discussed also apply to millimeter-wave radar–based human sensing [280], which has gained attention with the rise of commodity radars. Although mmWave radar uses different signal representations, such as range–Doppler and range–angle maps instead of Wi-Fi CSI, it faces similar challenges involving device heterogeneity, human body diversity, and environmental variability. Variations in radar hardware and antenna arrays parallel device heterogeneity in Wi-Fi; differences in body shape, motion, and orientation affect radar reflections similarly to CSI-based sensing; and environmental factors like clutter, multipath, and deployment geometry likewise influence robustness. Therefore, generalization-oriented strategies, e.g., multi-view sensing, large-scale in-the-wild datasets, domain adaptation, and continual learning, can be extended to mmWave radar sensing despite differences in signal modality.
@@ III Dataset
Recent years have witnessed a clear shift in Wi-Fi sensing research from lab-scale prototypes toward tools and datasets that support modern wireless standards and more complex, real-world sensing configurations. Early studies often relied on limited-bandwidth devices and small-scale, controlled environments, which constrained the evaluation of generalization performance. In contrast, emerging tools and benchmarks increasingly embrace higher bandwidths, denser subcarrier representations, and diverse deployment scenarios, enabling more realistic and robust system-level research.
Several open-source tools are available for extracting Wi-Fi Channel State Information and Beamforming Feedback Information (BFI), as summarized in Table VI. Among these, Linux CSI Tool [70] was the earliest open-source CSI extraction tool released in 2011, while PicoScenes [98, 121] represents a new generation of sensing tools that align with modern wireless standards. Specifically, PicoScenes supports advanced IEEE 802.11 standards up to Wi-Fi 7, offering bandwidths up to 320 MHz and up to 1024 subcarriers, which significantly expands the sensing resolution in both time and frequency domains and enables sensing under more complex multipath and heterogeneous deployment conditions. In addition, due to the limited availability of networks and devices supporting open CSI extraction, BFM-Tool [259] facilitates BFI extraction, enabling Wi-Fi sensing across a broader range of commercial devices.
TABLE VI: Wi-Fi Channel State Information and Beamforming Feedback Information extraction tools.
Name max. MIMO 802.11 support max. # subcarriers max. bandwidth Wi-Fi
Intel CSITool [70] $\times2$ n 60 40 MHz CSI
@@
Nexmon CSITool [65] $\times4$ n/ac 256 80 MHz CSI
Wi-ESP [7] $\times2$ b/g/n/ac 64 40 MHz CSI
[…]
PicoScenes [98, 121] $\times4$ a/g/n/ac/ax/be 1024 320 MHz CSI
BFM-Tool [259] $\times4$ ac/ax 512 160 MHz BFI
Beyond advances in data acquisition tools, recent Wi-Fi sensing datasets have also shifted from small-scale, lab-controlled collections toward large-scale, in-the-wild benchmarks. Over the years, many researchers have generously released public datasets, significantly improving the efficiency of algorithm evaluation by removing the need to collect data from scratch. Among the various sensing tasks, activity recognition is currently the most well-represented in public datasets, as summarized in Table VII. To provide a more intuitive overview of activity recognition datasets, Fig. 17 presents a bubble chart that visualizes the number of activity classes, release date, and dataset volume.
@@
Furthermore, to address the scarcity of data, promote industry-academia collaboration, and foster a thriving sensing ecosystem, we have initiated and built the Sensing Dataset Platform (SDP: https://www.sdp8.org/). As the most comprehensive open repository for wireless sensing datasets to date, SDP hosts a wide range of publicly available datasets contributed by both industrial and academic communities. The platform also encourages researchers to share their own data, aiming to foster open collaboration, accelerate innovation, and support reproducible research in the field of wireless sensing.
TABLE VII: Wi-Fi Datasets. The main results column mainly denotes the in-domain results from the original publications.
ID Dataset Main Domain Information Device Setting Highlight Main Results
Action Recognition
1 XRF55 [208] 4 environments, 39 participants, 55 actions 1Tx, 3Rx, 20MHz, 5GHz Multimodal, 55 actions 87.26%
2 BullyDetect [113] 8 environments, 20 paired participants, 7 actions 1Tx, 1Rx, 20MHz, 5GHz Bullying actions 90.4%
3 WiSDA [99] 1 environments, 3 participants, 6 actions 1Tx, 1Rx, 20MHz Widar3.0-like 99.5%
4 WiNDR [165] 1 environments, 3 participants, 5 actions, 24 orientation 1Tx, 1Rx, 20MHz, 5GHz Full 360° coverage 78%-92%
5 WiMANS [89] 3 environments, 6 participants, 9 actions 1Tx, 1Rx, 20MHz, 2.4/5GHz Multi-person 89.1%-96.6%
6 WiGuesture [309] 1 environments, 8 participants, 6 actions 1Tx, 1Rx, 100Hz, 2.4GHz ESP32S3 92.57%
7 RoGER [14] 2 environments, 4 participants, 6 actions, 4 orientation 1Tx, 2Rx, 20MHz, 2.437GHz AR9580 94.84%-99.71%
8 ImgFi [275] 1 environments, 5 participants, 6 actions 1Tx, 1Rx, 20MHz Converted to images 99.8%
9 ResMon [310] 2 environments, 3 areas, 6 participants, 4 actions 1Tx, 1Rx, 20/40MHz, 2.4/5GHz Cross-band test 83.80%-89.67%
10 Meneghello et al. [147] 6 environments, 4 participants, 7 actions 1Tx, 1Rx, 80MHz, 5.21GHz 80MHz not reported
11 Demrozi et al. [36] 2 environments, 6 participants, 6 actions 2Tx, 2Rx, 2.4GHz Nexmon firmware 99.3%
12 MM-Fi [254] 4 environments, 40 participants, 27actions 1TX, 1RX, 20MHz, 5GHz Multimodal not reported
13 NTU-Fi HAR [253] 1 environments, 20 participants, 6 actions 1Tx, 1Rx, 40MHz, 5GHz TP-Link N750 98.6%
14 FallDar [258] 4 environments, 6 locations, 6 participants, fall & normal 1Tx, 1Rx, 1000Hz, 5GHz 6 months 5.7% FAR
15 ReWis [9] 3 environments, 2 participants, 4 actions 1Tx, 3Rx, 20/80MHz, 5GHz 80MHz 78.25%-99.82%
16 OPERAnet [15] 2 environments, 6 participants, 6 actions 1Tx, 2Rx, 5GHz Multimodal 71%-100%
17 SHARP [148] 3 environments, 4 locations, 3 participants, 7 actions 1Tx, 1Rx, 80MB, 5GHz 80MHz 95.99%-99.79%
18 CSI-HAR-Dataset [151] 1 environments, 3 participants, 7 actions 1Tx, 1Rx, 20MHz, 5GHz 802.11ac 95.5%